i would also like help with this please
Write an equation in slope-intercept form that satisfies each condition.
slope -1/2,passes through (2,-3)
i did it like
-3=-1/2(2)-b
but i dont think thats right please help

Answer by checkley71(8403)

Y=mX+B OR IN THIS CASE WE HAVE Y=-X/2+b & -3=-2/2+b OR -3=-1+b OR b=-2 THUS
WE HAVE Y=-X/2-2
YOU WERE ON THE RIGHT TRACK.